OpenStudy (anonymous):

How do you calculate the energy of a photon with a frequency of 501 nm?

OpenStudy (anonymous):

It should be "a wavelength of 501nm." To calculate the energy of a photon, we use \(E=h\nu\) or \(E=\large\frac{hc}{\lambda}\). \(h\) is the Planck's constant, \(6.626\times 10 ^{-34}\) J.s \(c\) is the speed of light, usually \(3\times 10^8\) m/s \(\nu\) is the frequency, in Hz. \(\lambda\) is the wavelength, in m. In this problem, you are given the wavelength so we should use the second formula

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You are welcome